OpenStudy (anonymous):

Find three consecutive odd integers such that the sum of the largest and twice the smallest is 25. If x represents the smallest integer, then which equation could be used to solve the problem? 2x + x + 2 = 25 2x + x + 4 = 25 2x + 2x + 4 = 25 2x + 2x + 2 = 25

OpenStudy (anonymous):

numbers will be x,x+2,x+4

OpenStudy (anonymous):

2x + x + 2 = 25 2x + x + 4 = 25 2x + 2x + 4 = 25 2x + 2x + 2 = 25

OpenStudy (anonymous):

now according to the question (x+4)+2x=25

OpenStudy (anonymous):

2x + x + 2 = 25 2x + x + 4 = 25 2x + 2x + 4 = 25 2x + 2x + 2 = 25 One of the above...

OpenStudy (anonymous):

option 2